A 100 mL flask contained H2 at 200 Torr and a 200 mL flask contained He at 100 Torr. The two flask where then connected so that each gas filled their combined volume. Assuming no change in temperature, total pressur

Grade:12

1 Answers

Bhavya
7 years ago
Dear student
Considering T,n being constant
we can say that
P1V1 + P2V2 = PfVf (final, after two flasks are connected)
100*200 + 200*100 = Pf * 300
Pf = 400/3 torr